may/june 2017: 60Hz seen in alfa bm5,6 data.
17may17: investigating the problem
09jun17: resolving the problem.
17may17: Investigating the problem.
Project p2030 reported problems with alfa
beams 5 and 6 during observing runs in apr/may17. When processing
the runs, the rfi excision was excluding most of the beam 5 and
beam 6 data.
The p2030 setup is:
- alfa using both bands. one centered at 1450 MHz, the
other centered at 1300 MHz (each 172 MHz bw).
- parallactic angle tracking used.
- they typically look at a know pulsar (for 300 secs, no cal) at
- They then do 268 sec scans tracking each search position. The
hcorcal is fired for the first 5 seconds of each of these scans.
- Data is taken with the mock spectrometers:
- group 0 - psrfits files. 65 usec sampling 15270 spectra/row,
512 freq channels, polarization adding.
- group 1 - regular fits files, 1 second integrations for
Known problems with alfa beams:
- during the 07mar17 thermal cycling of the alfa dewar:
- beam1a which had been bad, started to work again
- beam5a which had been good, went bad.
- The lna's for alfa bm5a were turned off
- beam5a now has noise from the if/lo system
- We should split alfa 5b and inject it into alfa 5a and 5b
(making sure we have no reflections).
- This would improve the current signal to noise by sqrt(2).
This can be done downstairs in the if/lo output to the
Looking at the p2030 data.
I took a look at some of the p2030 data
that had been reported bad.
- beam 6, 1450 band. 512 frequency channels by 15270 spectra
(sampled at 65usecs)
- The dark lines are dropouts in power. They span the entire
frequency range (it was also seen in the 1300 MHz band).
- counting the dark lines, there are 12 in .2 secs .. or .01667
ms .. this is 60 Hz.
bm 6 total power plots vs time (10may17)
1450 MHz band (.ps) (.pdf)
10may17 1300 MHz band (.ps) (.pdf)
- This came from file:
- The total power was computed across the 172MHz band for each
65 usec sample
- The top frame plots 10 seconds of data.
- The jump at 6.3 secs is the hcorcal going off.
- The spectra was normalized to the cal off median.
- The bottom 2 frames plots .2 seconds around the cal
- The dashed green lines are spaced by 16.667 millisecs (60 Hz)
- middle frame: the plot is normalized to the cal on data
- The red lines show a 1 db step
- The bottom frame normalizes to the cal off median.
- The red lines show a 1 db step.
- The dropout last for about 6 millisecs.
- the falling edge is slow. The return is within 1 sample (65
- This looks like an attenuator stepping by 1db when it
- the only variable attenuator in the signal path is the 32 db
attenuator that we change in 1db steps (i think it might have a
resolution of .5db).
- actually i think there is also a 10db attenuator when
switching between narrow band (100 mhz) and wide band (300
- This is a pin diode attenuator, so it can probably change this
total power plots vs time from all 7 beams.
I looked at p2030 from apr17 thru may17 to see
what has been going on.
- for each date I looked at the start of a single source
- The total power for the first .2 seconds of the source was
plotted for all 7 beams.
- The data was normalized to the median value of the first
second of data
- Dashed red lines were plotted showing a 1 db step.
- The first page is the 1450 band, the 2nd page is the 1300 MHz
band (lots of radar rfi).
- 60 Hz jumps bms 5,6 group 0. < .5 db
- 60Hz jumps bms 5,6 group 0.
- 60Hz jumps bms 3,4,5,6 of group 1
- looks like pulsar search was taken with both
groups rather than grp 0 65usec, grp 1 1 sec..this
is an frb scan
- The jumps were not seen in bms 5,6 of group 0
(where they had been seen previously.
- 60Hz jumps bms 5,6 group 0 .. about .5 db
- 60Hz jumps bms 3,4,5,6 of group1
- looks similar to the 20apr17 data.
- 60 Hz jumps bms 5,6 group 0. < .5 db
- 60 Hz jumps bms 5,6 group 0. 1db jumps
- 60 Hz jumps bms 5,6 group 0 1 db jumps
- 60Hz jumps bms 5,6 group 0. .5 to 1db jumps
Taking data 16may17.
I took some data on 16may17 (after the p2030
- my setup was similar to p2030 (i used their config file).
- I did not see any 60 Hz jumps.
- i looked in the output data.
- We also looked at the output of the mock baseband mixers for
What is going on:
processing: 170515/alfabm56, alfabm56_more.pro
- 60 Hz usually means a power supply problem.
- When first looking at the 10may17 data with their 1db
steps, i thought we were looking at an attenuator problem.
- Looking at the earlier data, the jump was not always 1
db...So maybe it isn't an attenuator...
- but it could be that the fall time was slow enough that we
didn't get down to 1db before the 60Hz cycle caused it to jump
- The programmable attenuator is in the mock mixer
- for each pol of each beam
- there is an attenuator for the high and a separate
attenuator for the low band
- Each attenuator is driven by 5 bits coming from a micro
controller in the bbm chassis. (there is also a atten select
- There are two power supplies that drive the mixer chassis.
- 1 for beams 1..2
- 2 for beams 3..6
- luis looked at the ripple from these power supplies..
- He found that one of the power supplies did have a dead fan..
so maybe things are heating up..
- The fan will be replaced..
- We need to continue taking data until we can duplicate
170609: resolving the problem.
- On 7jun17 jason queried whether any progress had been made on
the alfa bm5,6 problem.
- Later that day i looked at alfa's output with a square law
detector/scope in the control room. This time i could see the 60
Hz dropouts. The problem was seen as the alfa signal entered the
control room.. so the problem was probably upstairs.
- Felix and crew went up and first inspected the rf to 250 Mhz
mixer chassis, but didn't find anything.
- After some more digging, it turned out that the problem was in
the bias cards of alfa.
- a 15 volt supply that fed beams 5 and 6 was bad.
- The bad supply was repaired, and replaced on 09jun17
After the repairs, i took some data with alfa to check things.
The setup was;
- center the 2 172 Mhz bands of alfa at 1300 and 1450 Mhz.
- use the mock spectrometers to take data:
- 172 Mhz bw on each band, 65 usec sampling, 16 bit spectra,
keep polA, B separate (p2030 adds the pols).
- take data for 60 seconds with the telescope stationary.
Processing the data:
- There were 918000 spectra each integrated for 65.3595 usecs.
- I only processed the 1450 Mhz band (the 1300 Mhz band has lots
- for each beam and pol I computed the total power for each
spectra, ending up with a total power time series of 918000.
- I normalized each time series to the median value (so
- y= abs(fft(totalPower))*2was computed and then plotted (*2
since i only kept 1/2 the points since a transform of real
the plots show the residual 60 hz
in the data (.ps) (.pdf)
- The top frame is polA, to bottom frame is polB
- Each beam is plotted in a separate color. I added an amplitude
offset to each beam for display.
- The beams were also shifted by bmNumber*1hz so that 60
hz from adjacent beams would not overlay each other.
- The amplitude of 60 Hz is about ,0005 to .0008 times Tsys..
- i checked this by processing a sine wave of unity amplitude
through the same code. It came out with an amplitude of 1.
- the 15 volt supply in the bias chassis feeding beams 5 and
6 had a 60 hz problem.
- It was replaced on 09jun17.
- after replacing it, the residual 60 hz through the mock
spectrometer was about .0006 Tsys.
- When first testing on 09jun16, i inserted the 100Mhz rf
filters in alfa.. to help with the rfi.
- I then saw strong 60 Hz on all beams of alfa.....
- so either the 100Mhz filter bank upstairs, or the extra amps
downstairs have a 60hz problem. Luckily no one uses the 100
Mhz filters anymore (they were initially installed when we
only had the wappsfor data taking.. with their 100Mhz
bandwidth limitation on alfa).